# Centre du Parc, Rue Marconi 19, CH-1920 Martigny, Switzerland

### v1

@@ Abstract
We present a novel approach to enhance the performance of sampling-based Model Predictive Control (MPC) in constrained optimization by leveraging products of experts. Our methodology divides the main problem into two components: one focused on optimality and the other on feasibility. By combining the solutions from each component, represented as distributions, we apply products of experts to implement a project-then-sample strategy. In this strategy, the optimality distribution is projected into the feasible area, allowing for more efficient sampling. This approach contrasts with the traditional sample-then-project method, leading to more diverse exploration and reducing the accumulation of samples on the boundaries. We demonstrate an effective implementation of this principle using a tensor train-based distribution model, which is characterized by its non-parametric nature, ease of combination with other distributions at the task level, and straightforward sampling technique. We adapt existing tensor train models to suit this purpose and validate the efficacy of our approach through experiments in various tasks, including obstacle avoidance, non-prehensile manipulation, and tasks involving staying on manifolds. Our experimental results demonstrate that the proposed method consistently outperforms known baselines, providing strong empirical support for its effectiveness.
@@ 1 Introduction
These methods have been widely used in motion planning (Orthey et al., 2023) and control problems (Bhardwaj et al., 2022). As they are typically gradient-free, they avoid the time needed for gradient computation, which is often significant. However, a primary challenge is the lack of efficient sampling mechanisms to capture essential information, particularly in cases involving constraints (especially equality constraints), sparse cost or reward functions, and high-dimensional problems. For instance, Fig. 1-a shows a robot trying to push a mustard bottle toward a target, but certain samples can lead it away from the object. Similarly, in Fig. 10, many samples fall outside the shaded area (the desired space for the robot’s motion), resulting in less informative or rejected samples. In this article, we address the challenge of efficient sampling in the presence of constraints by proposing a novel approach to enhance sampling efficiency.
@@
Figure 1: In sampling-based Model Predictive Control (SMPC), many samples can be discarded due to inefficiencies or infeasibility. In (a), only the green samples effectively influence the movement of the object, while the actions indicated by the red arrows have no impact on the environment. (b) illustrates the importance of selecting samples that do not collide with obstacles, which become more apparent as the robot approaches them; lighter colors indicate obstacles that appear later in the task.
Methods in the literature for constraint-based sampling can be categorized into three main approaches. The first category includes projection-based methods (Mirabel et al., 2016; Jinkyu Kim and Park, 2016), which sample first and then project into the feasible region. As shown in Fig. 3, even with an efficient projection mechanism (often difficult to implement), samples can accumulate near the boundary of the feasible region, leading the system to potentially converge toward that boundary. The second approach involves penalizing constraint violations (Bhardwaj et al., 2022), serving as a data-driven, iterative alternative to projection. However, this method does not guarantee feasible solutions before convergence and may require multiple iterations to achieve feasibility. The third approach, similar to Power and Berenson (2023) and Sacks and Boots (2023), learns a distribution that promotes sampling of feasible actions or states. While this approach can be effective, it is often complex to implement and computationally expensive (similarly to reinforcement learning methods).
To address these challenges and reduce the complexity of existing methods, we propose an approach that leverages multiple expert models, each specialized in a distinct aspect of the problem. Specifically, one expert is dedicated to learning the feasibility distribution, which focuses on constraints, while the other expert handles the optimality distribution. Moreover, as we show in this paper, the optimality distribution can be modeled as a Gaussian distribution using methods such as Model Predictive Path Integral (MPPI), so this part of the problem is computationally efficient and does not require additional learning, enabling rapid updates as the problem evolves.
Leveraging multiple expert models to simplify complex tasks is a widely adopted strategy in the machine-learning community. This is commonly implemented either through mixture of experts (MoE) (Jordan and Jacobs, 1993), which combines outputs in a manner analogous to an OR operation, or products of experts (PoE) (Hinton, 1999), which integrates outputs similarly to an AND operation. In our framework, since the objective is to identify solutions that are both optimal AND feasible, we adopt the products of experts paradigm. Beyond their inherent conceptual differences, PoE offers distinct advantages that make it particularly well-suited to our framework. Specifically, PoE enables each expert to focus on their own decision variables and objectives independently, simplifying and streamlining the learning process for each expert. For instance, in the example presented in this article, the expert tasked with optimizing for action optimality can concentrate solely on action variables (as in methods like MPPI), while another expert can focus on both action and state variables, or even exclusively on state variables, prioritizing feasibility without concern for optimization. By combining their outputs, PoE identifies the intersection of their respective distributions, yielding solutions that satisfy all variables and objectives simultaneously. In contrast, MoE typically combines solutions as a weighted average, potentially failing to account for such considerations.
Fig. 2 illustrates our proposed approach which splits the main problem into two sub-problems: the optimality problem, focusing on optimization, and the feasibility problem, focusing on constraints. We compute distributions for both subproblems and then combine them with the products of experts approach to create a distribution that is both feasible and optimal. This framework effectively projects the optimality distribution into the feasible region.
Unlike traditional projection methods that use a sample-then-project approach on individual samples (Stilman, 2010; Chavan-Dafle et al., 2020) or rely on tangent space approximations that require calculating system gradients (Pardo et al., 2015; Suh et al., 2011), our method employs a project-then-sample strategy, which increases the likelihood that samples fall within the feasible region and can be applied to systems with challenging-to-calculate gradients, which is for example the case for non-prehensile manipulation. This approach ensures that the feasibility distribution acts as a binary filter, effectively eliminating infeasible regions of the optimality distribution. Consequently, the success of our method hinges on accurately learning the feasibility distribution, which must precisely capture sharp transitions at boundaries.
In this article, we demonstrate how different experts can be effectively integrated using tensor train (TT) decomposition. Our previous work demonstrated the potential of this method for learning and sampling from desired distributions (Shetty et al., 2023; Xue et al., 2024a). We extend these findings by applying TT in the context of products of experts. We compare it to a baseline relying on normalizing flows (NFs) (Dinh et al., 2016), inspired by (Power and Berenson, 2023), to show the advantages of TT over NFs, as well as to discuss the differences between a project-then-sample strategy compared to a sample-then-project strategy.
In summary, our contributions are threefold:
  • Enhancing the sampling efficiency of SMPC methods by dividing the problem into multiple subproblems that address motion feasibility and optimality and combining them using products of experts (i.e., project-then-sample strategy). This enables more effective exploration of the feasible set and extends applicability to problems where gradient calculation is challenging.
  • Introducing an effective approach to leverage products of experts through TT density estimation. In particular, we build upon (Shetty et al., 2023) to show how this method integrates with other SMPC approaches for longer horizon planning.
@@ 2 Related work
Sampling-based MPC (SMPC): Methods such as Rapidly-Exploring Random Tree (RRT) (Kuffner and LaValle, 2000) and Probabilistic Roadmap Planner (PRM) (Kavraki et al., 1996) for motion planning, as well as approaches such as Cross-Entropy Motion Planning (CEM) (Kobilarov, 2012) and Model Predictive Path Integral (MPPI) (Williams et al., 2017b) have shown the effectiveness of sampling strategies in addressing MPC problems. Recent findings by DeepMind (Howell et al., 2022) suggest that even for a challenging task such as quadruped walking, selecting the best sample suffices to achieve the objective. However, these advantages are limited to tasks without substantial constraints.
@@
Constrained Sampling-based Motion Planning: Given that sampling-based methods offer probabilistic completeness, they may require extensive data, posing challenges to finding constraint-free solutions for highly constrained problems, especially in higher dimensions with equality constraints. Projection methods are widely used in motion planning, particularly within RRT, to ensure constraints are met and improve sampling efficiency (Kingston et al., 2018). Prior research shows that this approach enables probabilistic completeness (Berenson, 2011), ensuring that the feasible set will be fully explored over time. However, finding an analytical solution for projection remains challenging under most constraints. In specific cases, projection operators can be locally defined using the tangent space and a nullspace projector (Pardo et al., 2015) or by direct sampling within the tangent space (Suh et al., 2011). Nevertheless, this approach has limitations, as it only captures local constraint behavior and fails with non-differentiable constraints (e.g., contact-rich manipulation). To address this, sampling-based methods often use gradient-based techniques to approximate the projection (Stilman, 2010; Yakey et al., 2001). Some works, such as (Stilman, 2010), frame projection as a secondary optimization problem. An alternative is to relax constraints by introducing penalty terms within the main objective function (Bhardwaj et al., 2022). However, both approaches face limitations, including increased computational demands and the risk of converging to infeasible regions.
@@
This article contends that the above approaches may not be the most efficient, even with an analytical projection solution. Standard projection-based sampling may guide exploration toward regions with limited feasibility, yielding biased solutions and insufficiently exploring the feasible set. A more effective approach could be to directly model the distribution of the feasible set and sample exclusively from this distribution, concentrating exploration solely within feasible regions. Recent advances in generative models have facilitated this, allowing feasibility-aware sampling strategies (Sacks and Boots, 2023; Power and Berenson, 2023). However, these methods, akin to RL, often require learning highly complex distributions, which can be data-intensive and computationally costly.
@@
Product of Experts: The utilization of the product of experts was initially introduced by Hinton (1999), wherein each expert represents a relatively simple distribution focusing on specific parameters or tasks. These experts are subsequently combined to yield a more refined and sharper distribution. This methodology has found applications in robotics, including the fusion of sensory information (Pradalier et al., 2003), the fusion of robot motion learned in multiple coordinate systems (Calinon, 2016), and across task spaces (Pignat et al., 2022; Mühlbauer et al., 2024). Our work extends these prior approaches by incorporating the product of experts framework within an MPC paradigm. One of the key challenges we address in this article is that some distributions (e.g., feasibility) are often too intricate to be accurately captured by simple Gaussian models. Previously mentioned works in PoE frameworks have relied on Gaussian representations due to their mathematical tractability. Gaussian distributions simplify product calculations and facilitate efficient sampling because analytical solutions for these operations are readily available. However, when dealing with highly non-linear or discontinuous feasibility constraints, Gaussian models fall short in capturing the sharp transitions and boundaries that characterize such distributions. If alternative methods were employed, additional steps might be required to normalize the resulting distribution and to develop effective sampling techniques, both of which would entail significant computational and implementation challenges. We demonstrate that with our proposed approach, these two steps become unnecessary.
@@ 3.1 Model Predictive Path Integral
The primary aim of this methodology is to tackle an optimal control problem associated with a dynamic system described by $\bm{x}_{t+1}=f(\bm{x}_{t},\bm{u}_{t},\bm{\omega}_{t})$ within a receding horizon $H$. Here, $\bm{x}_{t}\in\mathbb{R}^{d_{x}}$ and $\bm{u}_{t}\in\mathbb{R}^{d_{u}}$ denote the system state and action at time $t$, respectively, while $\bm{\omega}_{t}$ represents a stochastic noise element in the system. This objective is achieved by minimizing a predetermined cost function as
$c(\hat{\bm{X}}_{t},\hat{\bm{U}}_{t})=\sum_{h=0}^{H-1}c_{t}(\hat{\bm{x}}_{t+h},\hat{\bm{u}}_{t+h})+c_{H}(\hat{\bm{x}}_{t+H}).$ (1)
Here, $\hat{\bm{X}}_{t}=[\hat{\bm{x}}_{t}^{\scriptscriptstyle\top},\hat{\bm{x}}_{t+1}^{\scriptscriptstyle\top},\ldots,\hat{\bm{x}}_{t+H}^{\scriptscriptstyle\top}]^{\scriptscriptstyle\top}$ represents the state trajectory, and $\hat{\bm{U}}_{t}=[\hat{\bm{u}}_{t}^{\scriptscriptstyle\top},\hat{\bm{u}}_{t+1}^{\scriptscriptstyle\top},\ldots,\hat{\bm{u}}_{t+H-1}^{\scriptscriptstyle\top}]^{\scriptscriptstyle\top}$ denotes the sequence of action commands. Additionally, $c_{t}$ and $c_{H}$ denote stage and terminal cost values, respectively. Here, the constraints are formulated inside the cost function. The objective is to determine a set of optimal action commands $\bm{U}_{t}^{*}$ that minimizes this cost function while taking into account the dynamics of the system, specifically
@@
Williams et al. (Williams et al., 2017a, b) demonstrated that the solution for this optimal control problem can be effectively approximated using a forward approach, diverging from the conventional dynamic programming method. This discovery opened avenues for leveraging Monte Carlo estimation in approximating optimal actions through sampling-based methods. Furthermore, Wagener et al. (2019) showed that if the system policy can be modeled as a Gaussian distribution such as
@@
then the policy parameters can be approximated as
$\bm{\mu}_{t+h}=(1-\gamma_{t})\hat{\bm{\mu}}_{t+h}+\gamma_{t}\sum_{i=1}^{N}w_{i}\hat{\bm{u}}_{t+h},$ (4)
where $N$ denotes the number of samples, $\hat{\bm{\mu}}_{t+h}$ represents the preceding mean value, and $\gamma_{t}$ is the step size. The term $w_{i}$ is a weighting factor expressed as
@@
where $\beta$ is a temperature parameter. Following the adjustment of the policy parameters, the initial action $\bm{u}^{*}_{t}$ can be sampled, applied to the system, and iteratively repeated until the task is successfully completed.

### v2

@@ Abstract
We present a novel approach to enhance the performance of sampling-based Model Predictive Control (MPC) in constrained optimization by leveraging products of experts. Our methodology divides the main problem into two components: one focused on optimality and the other on feasibility. By combining the solutions from each component, represented as distributions, we apply products of experts to implement a project-then-sample strategy. In this strategy, the optimality distribution is projected into the feasible area, allowing for more efficient sampling. This approach contrasts with the traditional sample-then-project and naïve sample-then-reject method, leading to more diverse exploration and reducing the accumulation of samples on the boundaries. We demonstrate an effective implementation of this principle using a tensor train-based distribution model, which is characterized by its non-parametric nature, ease of combination with other distributions at the task level, and straightforward sampling technique. We adapt existing tensor train models to suit this purpose and validate the efficacy of our approach through experiments in various tasks, including obstacle avoidance, non-prehensile manipulation, and tasks involving staying in a restricted volume. Our experimental results demonstrate that the proposed method consistently outperforms known baselines, providing strong empirical support for its effectiveness. Sample codes for this project are available at https://github.com/idiap/smpc_poe
@@ 1 Introduction
For a robot to navigate and interact effectively within its environment, it must plan its motion while taking surrounding conditions into account. In most cases, the robot must satisfy various constraints imposed by physics, the environment, or its own mechanical limitations. Adhering to these constraints is often what makes motion planning and control a challenging problem in robotics.
@@
(a) Planar Pushing
(b) Whole-body Obstacle Avoidance
@@
Figure 1: In sampling-based Model Predictive Control (SMPC), many samples can be discarded due to inefficiencies or infeasibility. In (a), only the green samples effectively influence the movement of the object, while the actions indicated by the red arrows have no impact on the environment. (b) illustrates the importance of selecting samples that do not collide with obstacles.
These methods have been widely applied in trajectory optimization (Orthey et al., 2023), including well-known techniques like Rapidly-Exploring Random Trees (RRT) (Kuffner and LaValle, 2000). Their gradient-free nature makes them suitable for applications where computing gradients is difficult or infeasible. However, as these methods are only probabilistically complete, they may require extensive sampling. To address this, a common strategy is to reduce the planning horizon and focus on local decision-making. In this context, sampling-based model predictive (SMPC) approaches such as (Williams et al., 2017b; Bhardwaj et al., 2022) have been developed to improve responsiveness and adaptability. However, a primary challenge is the lack of efficient sampling mechanisms to capture essential information, particularly in cases involving constraints, sparse cost or reward functions. For instance, Fig. 1-a shows a robot trying to push a mustard bottle toward a target, but certain samples can lead it away from the object. Similarly, in Fig. 1-b, many samples can collide with the obstacles, resulting in less informative or rejected samples. In this article, we address the challenge of efficient sampling in the presence of constraints by proposing a novel approach to enhance sampling efficiency.
@@
A basic approach to incorporating constraints into sampling-based methods is to generate samples from a distribution naively without considering constraints and subsequently reject those that violate the constraints. This sample-then-reject strategy can be highly inefficient in tightly constrained problems, leading to a large number of discarded samples. This inefficiency has motivated the development of numerous alternative approaches in the literature that improve the sampling process to reduce the number of rejected samples. These methods can be categorized into three main approaches. The first category includes projection-based methods (Mirabel et al., 2016; Jinkyu Kim and Park, 2016), which sample first and then project into the feasible region. However, projection is not always straightforward: some methods (Pardo et al., 2015; Suh et al., 2011) require iterative computation of system gradients to perform the projection, which can be problematic when gradients are difficult to obtain—such as in non-prehensile manipulation. As shown in Fig. 2, even with an efficient projection mechanism (often difficult to implement), samples can accumulate near the boundary of the feasible region, leading the system to potentially converge toward that boundary. The second approach involves penalizing constraint violations (Bhardwaj et al., 2022), serving as a data-driven, iterative alternative to projection. However, this method does not guarantee feasible solutions before convergence and may require multiple iterations to achieve feasibility. The third approach, similar to Power and Berenson (2023) and Sacks and Boots (2023), learns a distribution that promotes sampling of feasible and optimal actions or states. While this approach can be effective, it is often complex to implement and computationally expensive (similarly to reinforcement learning methods).
One idea to solve these complexities can be to use several experts. Leveraging multiple expert models to simplify complex tasks is a widely adopted strategy in the machine-learning community. This is commonly implemented either through mixture of experts (MoE) (Jordan and Jacobs, 1993), which combines outputs in a manner analogous to an OR operation, or products of experts (PoE) (Hinton, 1999), which integrates outputs similarly to an AND operation. In our framework, since the objective is to identify solutions that are both optimal AND feasible, we adopt the products of experts paradigm. Beyond their inherent conceptual differences, PoE offers distinct advantages that make it particularly well-suited to our framework. Specifically, PoE enables each expert to focus on their own decision variables and objectives independently, simplifying and streamlining the learning process for each expert. For instance, in the example presented in this article, the expert tasked with optimizing for action optimality can concentrate solely on action variables (as in methods like MPPI), while another expert can focus on both action and state variables, or even exclusively on state variables, prioritizing feasibility without concern for optimization. By combining their outputs, PoE identifies the intersection of their respective distributions, yielding solutions that satisfy all variables and objectives simultaneously. In contrast, MoE typically combines solutions as a weighted average, potentially failing to account for such considerations.
Similarly, in other works on SMPC, such as (Pan et al., 2024), researchers have employed multiple experts to define the overall distribution. In these approaches, each expert specifies a distribution over the full trajectory, and the experts are combined first, either through a summation of cost terms or a product of distributions, before jointly solving the resulting problem. In contrast, our method decomposes the problem differently. The feasibility expert is learned independently and then combined with the optimality distribution. By taking the product of these two distributions, we reduce the probability of the optimality distribution in the infeasible area, thereby improving the efficiency of the sampling process. The optimality distribution can be modeled as a Gaussian distribution using methods such as Model Predictive Path Integral (MPPI), so this part of the problem is computationally efficient and does not require any learning. By introducing a feasibility expert and integrating it with optimality considerations, we obtain the project-then-sample strategy, which efficiently focuses the search on the feasible region of the optimization problem, thereby improving overall efficiency.
Fig. 3 illustrates our proposed approach, which splits the main problem into two sub-problems: the optimality problem, focusing on optimization, and the feasibility problem, focusing on constraints. We compute distributions for both subproblems and then combine them with the products of experts approach to create a distribution that is both feasible and optimal. This framework effectively projects the optimality distribution into the feasible region. A key remaining challenge is how to learn these typically complex, multimodal distributions and to sample from them efficiently.
There are three main paradigms for combining experts at the solution level in the literature: Gaussian experts (Calinon, 2016; Hinton, 1999), energy-based models (Gkanatsios et al., 2023), and diffusion-based models (Yang et al., 2023; Wang et al., 2024). Gaussian experts offer closed-form inference but are inherently restrictive in the kinds of multimodal distributions they can represent. Energy-based models are more flexible but impose heavy sampling burdens (e.g., Langevin dynamics or MCMC). Diffusion-based approaches significantly reduce sampling complexity, yet typically require large demonstration datasets. Critically, these models often fail to consider feasibility when they are performing sampling during the denoising process, leading to the rejection of many samples at each step. In this paper, we show that by integrating a dedicated feasibility expert, we can improve its reliability before sampling and boost per-step sampling efficiency throughout the diffusion process.
In this article, we demonstrate how different experts can be effectively integrated using tensor train (TT) decomposition. Our previous work demonstrated the potential of this method for learning and sampling from desired distributions (Shetty et al., 2024a; Xue et al., 2024a). We extend these findings by applying TT in the context of products of experts. We compare it to a baseline relying on normalizing flows (NFs) (Dinh et al., 2016), inspired by (Power and Berenson, 2023), to show the advantages of TT over NFs, as well as to discuss the differences between a project-then-sample strategy compared to a sample-then-project or pure sample-then-reject strategies.
In summary, our contributions are threefold:
  • Enhancing the sampling efficiency of SMPC methods by introducing the feasibility expert and combining it with the optimality distribution using products of experts (i.e., project-then-sample strategy). Particularly, the feasibility distribution is learned separately and then combined with the optimality one before sampling. This enables more effective exploration of the feasible set and extends applicability to problems where gradient calculation is challenging.
  • Introducing an effective approach to leverage products of experts through TT density estimation, a non-parametric distribution capable of learning multi-modal distributions. This approach can also leverage existing straightforward sampling algorithms, enabling efficient exploration of complex distribution spaces.
@@ 2 Related work
Gradient-Based vs Sampling-based MPC: Analytical solutions are often impractical for optimization problems in robotics, especially with non-convex, non-linear cost functions and constraints arising from kinematic chains and dynamics. Gradient-based methods like iLQR (Todorov and Jordan, 2002; Diehl et al., 2005) are widely used, offering scalability and local solution updates. However, they face issues such as local minima, and undefined gradients in tasks like contact-rich manipulation (Pang et al., 2023; Dafle et al., 2014). These limitations are critical in Model Predictive Control (MPC), where fast solutions are essential.
To address these challenges, sampling-based methods have gained attention. They bypass the need for gradients, can exploit GPU-based simulators for parallel sampling (Liang et al., 2018), and may reduce sensitivity to local minima by approximating smoothed gradients (Pang et al., 2023). This makes them promising alternatives for complex robotic tasks involving discrete decisions (Chavan-Dafle et al., 2020).
@@
Sampling-based MPC (SMPC): Methods such as Rapidly-Exploring Random Tree (RRT) (Kuffner and LaValle, 2000) and Probabilistic Roadmap Planner (PRM) (Kavraki et al., 1996) for trajectory optimization, as well as approaches such as Cross-Entropy Motion Planning (CEM) (Kobilarov, 2012) and Model Predictive Path Integral (MPPI) (Williams et al., 2017b) have shown the effectiveness of sampling strategies in addressing MPC problems. Recent findings by DeepMind (Howell et al., 2022) suggest that even for a challenging task such as quadruped walking, selecting the best sample suffices to achieve the objective. However, these advantages are limited to tasks without substantial constraints.
@@
Constrained Sampling-based Motion Planning: Given that sampling-based methods offer probabilistic completeness, they may require extensive data, posing challenges to finding constraint-free solutions for highly constrained problems. Projection methods are widely used in sampling-based methods, particularly within RRT, to ensure constraints are met and improve sampling efficiency (Kingston et al., 2018). Prior research shows that this approach enables probabilistic completeness (Berenson, 2011), ensuring that the feasible set will be fully explored over time. However, finding an analytical solution for projection remains challenging under most constraints. In specific cases, projection operators can be locally defined using the tangent space and a nullspace projector (Pardo et al., 2015) or by direct sampling within the tangent space (Suh et al., 2011). Nevertheless, this approach has limitations, as it only captures local constraint behavior and fails with non-differentiable constraints (e.g., contact-rich manipulation). To address this, sampling-based methods often use gradient-based techniques to approximate the projection (Stilman, 2010; Yakey et al., 2001). Some works, such as (Stilman, 2010), frame projection as a secondary optimization problem. An alternative is to relax constraints by introducing penalty terms within the main objective function (Bhardwaj et al., 2022). However, both approaches face limitations, including increased computational demands and the risk of converging to infeasible regions.
@@
This article contends that the above approaches may not be the most efficient, even with an analytical projection solution. Standard projection-based sampling may guide exploration toward regions with limited feasibility (e.g., boundaries), yielding biased solutions and insufficiently exploring the feasible set. A more effective approach could be to directly model the distribution of the feasible and optimal set and sample exclusively from this distribution, concentrating exploration solely within feasible regions. Recent advances in generative models have facilitated this, allowing feasibility-aware sampling strategies (Sacks and Boots, 2023; Power and Berenson, 2023). However, these methods, akin to RL, often require learning highly complex distributions, which can be data-intensive and computationally costly.
@@
Product of Experts: In SMPC, it is common practice to define the overall cost function or solution distribution by combining multiple terms, typically through summation or a product of distributions (Pan et al., 2024). Each term is designed to address a specific aspect of the problem, such as optimality or constraint satisfaction, and the final solution is obtained by solving the joint problem defined by their combination.
In most existing approaches, this combination occurs at the definition level, meaning that the full problem is formed first and then solved using techniques such as gradient descent, Langevin dynamics, or other sampling-based methods. While this is effective, it tends to reduce flexibility, as the individual structures of the constituent terms are lost once they are merged.
In contrast, Hinton’s original formulation of the product of experts (Hinton, 1999) suggests an alternative: solving each expert individually and then combining their solutions. By solution, we refer to a sampleable distribution, such as a Gaussian, that does not require further optimization or inference steps. This approach preserves the modularity of each expert, enables the use of specialized solvers, and simplifies sampling from the final distribution without re-solving the full joint problem. In this regard, each expert represents a relatively simple distribution focusing on specific parameters or tasks. These experts are subsequently combined to yield a more refined and sharper distribution. This methodology has found applications in robotics, including the fusion of sensory information (Pradalier et al., 2003), the fusion of robot motion learned in multiple coordinate systems (Calinon, 2016), and across task spaces (Pignat et al., 2022; Mühlbauer et al., 2024). Our work extends these prior approaches by incorporating the product of experts framework within an MPC paradigm.
There are three main approaches in the literature for combining multiple experts at the solution level: (1) Gaussian experts (Calinon, 2016; Hinton, 1999), (2) Energy-based models (Gkanatsios et al., 2023), and (3) Diffusion-based models (Yang et al., 2023; Wang et al., 2024). Each of these methods offers different advantages and trade-offs, which we discuss further to motivate our use of the TT representation.
Compared to these methods, Gaussian distributions are significantly more limited than TT and other baselines, primarily due to the parametric nature of Gaussian models. Energy-based models (LeCun et al., 2006), which can be viewed as a more general form of Gaussian distributions, indeed hold greater potential. However, according to our experience, sampling from these models typically relies on Langevin dynamics or methods such as MCMC, which can become a significant computational bottleneck. In contrast, TT does not suffer from this issue and allows for efficient and straightforward sampling.
Diffusion models (Ho et al., 2020) (or similar variations such as Flow-Matching (Lipman et al., 2023)) address the sampling challenges present in energy-based models to a significant extent. These models can be broadly categorized into two types: data-based (Wang et al., 2024) and model-based (Pan et al., 2024). Data-based methods require the optimality distribution to be explicitly learned in diffusion form, which typically demands a substantial number of high-quality demonstrations—a requirement not assumed in our work.
Moreover, recent works such as (Xue et al., 2025; Pan et al., 2024) have proposed innovative strategies to represent diffusion models in a data-free manner. Their approach, which bears resemblance to annealed MPPI methods, combines multiple experts at the function level. In their pipeline, different candidate solutions are sampled first, followed by a rejection step to discard infeasible samples at each denoising step. By contrast, our approach merges experts at the solution level. This allows us to reshape the sampling distribution before sampling, which is a more proactive method than relying solely on post-hoc sample rejection. Concretely, we inject the feasibility expert’s density into every denoising (sampling) step of the diffusion process, producing candidates that are inherently more likely to satisfy constraints.
@@ 3.1 Model Predictive Path Integral
The primary aim of this methodology is to tackle an optimal control problem associated with a dynamic system described by $\bm{x}_{t+1}=f_{\text{dyn}}(\bm{x}_{t},\bm{u}_{t},\bm{\omega}_{t})$ within a receding horizon $H$. Here, $\bm{x}_{t}\in\mathbb{R}^{d_{x}}$ and $\bm{u}_{t}\in\mathbb{R}^{d_{u}}$ denote the system state and action at time $t$, respectively, while $\bm{\omega}_{t}$ represents a stochastic noise element in the system. This objective is achieved by minimizing a predetermined cost function as
$c(\hat{\bm{X}}_{t},\hat{\bm{U}}_{t})=\sum_{h=0}^{H-1}c_{h}(\hat{\bm{x}}_{t+h},\hat{\bm{u}}_{t+h})+c_{H}(\hat{\bm{x}}_{t+H}).$ (1)
Here, $\hat{\bm{X}}_{t}=[\hat{\bm{x}}_{t}^{\scriptscriptstyle\top},\hat{\bm{x}}_{t+1}^{\scriptscriptstyle\top},\ldots,\hat{\bm{x}}_{t+H}^{\scriptscriptstyle\top}]^{\scriptscriptstyle\top}$ represents the state trajectory, and $\hat{\bm{U}}_{t}=[\hat{\bm{u}}_{t}^{\scriptscriptstyle\top},\hat{\bm{u}}_{t+1}^{\scriptscriptstyle\top},\ldots,\hat{\bm{u}}_{t+H-1}^{\scriptscriptstyle\top}]^{\scriptscriptstyle\top}$ denotes the sequence of action commands. The hat symbol $\hat{(\cdot)}$ indicates that these actions are part of the planned, but not yet applied, control sequence. Additionally, $c_{h}$ and $c_{H}$ denote stage and terminal cost values, respectively. Here, the constraints are formulated inside the cost function. The objective is to determine a set of optimal action commands $\bm{U}_{t}^{*}$ that minimizes this cost function while taking into account the dynamics of the system, specifically
@@
Williams et al. (Williams et al., 2017a, b) demonstrated that the solution for this optimal control problem can be effectively approximated using a forward approach, diverging from the conventional dynamic programming method. This discovery opened avenues for leveraging Monte Carlo estimation in approximating optimal actions through sampling-based methods. Furthermore, Wagener et al. (2019) showed that if the system motion can be modeled as a Gaussian distribution such as
@@
then its parameters can be approximated as
$\bm{\mu}_{t}=(1-\gamma_{t})\hat{\bm{\mu}}_{t}+\gamma_{t}\sum_{i=1}^{N}w_{i}\hat{\bm{U}}^{(i)}_{t}$, (4)
where $N$ denotes the number of samples, $\hat{\bm{\mu}}_{t}$ represents the preceding mean value, and $\gamma_{t}$ is the step size. The term $w_{i}$ is a weighting factor expressed as
@@
where $\beta$ is a temperature parameter. Following the calculation of the parameters, the initial action $\bm{u}^{*}_{t}$ can be sampled, applied to the system, and iteratively repeated until the task is successfully completed.